OpenStudy (pumpkin1102):

Jill wants to solve the following system using the elimination method: y = x + 8 y = 3x − 11 What number should the equation y = x + 8 be multiplied by to eliminate x? −1 −2 −3 3

OpenStudy (welshfella):

you need to chANGE THE X IN THE FIRST EQUATION TO -3X

OpenStudy (welshfella):

to do that you multiply the first equation by what?

OpenStudy (pumpkin1102):

idk I am so confused on this

OpenStudy (welshfella):

x times ___ = -3x what goes in the blank space?

OpenStudy (pumpkin1102):

y?

OpenStudy (welshfella):

no that would give you xy

OpenStudy (welshfella):

you need -3x

OpenStudy (pumpkin1102):

I have no idea

OpenStudy (welshfella):

x times 2 = 2x x times 4 = 4x x times -3 = -3x

OpenStudy (welshfella):

simple

OpenStudy (pumpkin1102):

so basically the number without x?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

y = x + 8 * -3 is -3y = -3x - 24 when you add this equation to y = 3x - 11 the terms in x disappear because -3x + 3x = 0

OpenStudy (pumpkin1102):

thx

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

terms like -3x mean '-3 times x'.

OpenStudy (welshfella):

the multiply is understood

OpenStudy (pumpkin1102):

angt iu went ih the ohter?

OpenStudy (pumpkin1102):

wiw gitch

OpenStudy (welshfella):

???